OpenStudy (anonymous):

2nd derivative i have the first!

OpenStudy (anonymous):

find the derivative of the first derivative

OpenStudy (anonymous):

original is \[x \sqrt{9-x}\]first is \[\frac{ -3x+18 }{ 2\sqrt{9-x} }\]

OpenStudy (anonymous):

then you use the quotient rule

OpenStudy (anonymous):

i just need help simplifying

OpenStudy (anonymous):

where i'm at

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

yea, just try to find out the easiest way to type = ="

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3(x-12)/(4(9-x)^(3/2))

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

first, factor out constants: 1/2(d/dx(18-3x)/sqrt(9-x)) use the product rule: and you will get \[\frac{ 1 }{ 2 }\left[(18-3x)( d/dx\frac{ 1 }{\sqrt{9-x} }) \right]+(d/dx\frac{ 1 }{ \sqrt{9-x} })\] Differentiate the sum term by term and factor out constant: (first attach file) Using the chain rule, and you will get (2nd attach pic) derivative of 18 is zero, and so you will get (3rd) differentiate the sum term by term and factor out constant (4th) the derivative of x is 1 (5th) the derivative of 9 = 0 (6th) finally the derivative of x =1 (7th)

OpenStudy (anonymous):

actually, never mind what attachment, each shows you complete solution